OpenStudy (anonymous):

The track team has 11 distance runners but can only enter five in the meet. How many combinations of five runners can the coach create? 462 120 55,440 3,696

OpenStudy (anonymous):

Please help.

OpenStudy (kropot72):

\[nCr=\frac{n!}{(n-r)!r!}=\frac{11!}{6!5!}\] \[11C5=\frac{11\times 10\times 9\times 8\times7}{5\times 4\times 3\times 2}\]